# Modeling Binary Lenses and Sources with the BAGLE Python Package

### v1

@@ Appendix B $u_{0}$ transformation
$\boldsymbol{u}_{0}^{\prime}$ $=$ $S^{\prime R^{\prime}}-L^{\prime R^{\prime}}=(S^{\prime R}-d\hat{s})-(L^{\prime R}-d\hat{s})=S^{\prime R}-L^{\prime R}$ (B13)
$\boldsymbol{u}_{0}^{\prime}$ $=$ $\boldsymbol{u}_{0}+\frac{\boldsymbol{\mu}_{rel}}{\theta_{E}}(t_{0}^{\prime}-t_{0})-d\hat{s}$ (B14)
@@
$\boldsymbol{\mu}_{rel}$ $=$ $[\mu_{rel}\cos\theta_{\mu},\mu_{rel}\sin\theta_{\mu}]$ (B15)
$\hat{s}$ $=$ $[\cos\theta_{s},\sin\theta_{s}]$ (B16)
Where $\theta_{\mu}$ is the angle from North to $\boldsymbol{\mu}_{rel}$ and $\theta_{s}$ is the angle from North to the binary axis, East of North. (Note that choosing North as our reference will not affect the final answer. Another reference could be chosen). So in components:
$u_{0,E}^{\prime}$ $=$ $u_{0,E}+\mu_{rel}\cos\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\cos\theta_{s}$ (B17)
$u_{0,N}^{\prime}$ $=$ $u_{0,N}+\mu_{rel}\sin\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\sin\theta_{s}$ (B18)
@@
$C\equiv(\hat{\mu}_{rel}\times\hat{u}_{0})\cdot\hat{z}$ (B20)
@@
$\theta_{u}=\theta_{\mu}-90^{\circ}C$ (B21)
We can now take Eq. B14 and define the components as their magnitudes times cos/sin of angles:
$u_{0}^{\prime}\cos\theta_{u^{\prime}}$ $=$ $u_{0}\cos\theta_{u}+\mu_{rel}\cos\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\cos\theta_{s}$ (B22)
$u_{0}^{\prime}\sin\theta_{u^{\prime}}$ $=$ $u_{0}\sin\theta_{u}+\mu_{rel}\sin\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\sin\theta_{s}$ (B23)
Where $\theta_{u^{\prime}}$ is the angle from North to $u_{0}^{\prime}$. $u_{0}$ will always be parallel to $u_{0}^{\prime}$, but they may be opposite directions. So
$\theta_{u^{\prime}}=\theta_{u}+180^{\circ}F$ (B24)
@@
$\sin\theta_{u^{\prime}}$ $=$ $G\sin\theta_{u}$ (B25)
$\cos\theta_{u^{\prime}}$ $=$ $G\cos\theta_{u}$ (B26)
where G = -1 if F = 1 and G = 1 if F = 0. Plugging into Eqs B22 and B23:
$u_{0}^{\prime}G-u_{0}$ $=$ $\frac{1}{\cos\theta_{u}}\Big(\mu_{rel}\cos\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\cos\theta_{s}\Big)$ (B27)
$u_{0}^{\prime}G-u_{0}$ $=$ $\frac{1}{\sin\theta_{u}}\Big(\mu_{rel}\sin\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\sin\theta_{s}\Big)$ (B28)
@@
$\cos\theta_{u}$ $=$ $\cos(\theta_{\mu}-90^{\circ}C)=C\sin\theta_{\mu}$ (B29)
$\sin\theta_{u}$ $=$ $\sin(\theta_{\mu}-90^{\circ}C)=-C\cos\theta_{\mu}$ (B30)
@@
$u_{0}^{\prime}G-u_{0}$ $=$ $\frac{1}{C}\Big(\mu_{rel}\frac{1}{\tan\theta_{\mu}}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\frac{\cos\theta_{s}}{\sin\theta_{\mu}}\Big)$ (B31)
$u_{0}^{\prime}G-u_{0}$ $=$ $-\frac{1}{C}\Big(\mu_{rel}\tan\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\frac{\sin\theta_{s}}{\cos\theta_{\mu}}\Big)$ (B32)
@@
$\sin\theta_{s}$ $=$ $\sin(\theta_{\mu}+\phi)=\cos\phi\sin\theta_{\mu}+\cos\theta_{\mu}\sin\phi$ (B33)
$\cos\theta_{s}$ $=$ $\cos(\theta_{\mu}+\phi)=\cos\phi\cos\theta_{\mu}-\sin\phi\sin\theta_{\mu}$ (B34)
$\rightarrow\frac{\sin\theta_{s}}{\cos\theta_{\mu}}$ $=$ $\cos\phi\tan\theta_{\mu}+\sin\phi$ (B35)
$\frac{\cos\theta_{s}}{\sin\theta_{\mu}}$ $=$ $\frac{\cos\phi}{\tan\theta_{\mu}}-\sin\phi$ (B36)
$\rightarrow\frac{\sin\theta_{s}}{\cos\theta_{\mu}}+\frac{\cos\theta_{s}}{\sin\theta_{\mu}}$ $=$ $\cos\phi\Big(\frac{1}{\tan\theta_{\mu}}+\tan\theta_{\mu}\Big),$ (B37)
Setting Eq B31 equal to Eq B32 and simplifying:
$\mu_{rel}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)\Big(\frac{1}{\tan\theta_{\mu}}+\tan\theta_{\mu}\Big)$ $=$ $d\Big(\frac{\sin\theta_{s}}{\cos\theta_{\mu}}+\frac{\cos\theta_{s}}{\sin\theta_{\mu}}\Big)$ (B38)
$\mu_{rel}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)$ $=$ $d\cos\phi$ (B39)
We can now use this relation in Eqs B17 and B18. Starting with the E-component:
$u_{0,E}^{\prime}$ $=$ $u_{0,E}+d\cos\theta_{\mu}\cos\phi-d\cos\theta_{s}$ (B40)
$u_{0,E}^{\prime}$ $=$ $u_{0,E}+d(\cos\theta_{\mu}\cos\phi-(\cos\phi\cos\theta_{\mu}-\sin\theta_{\mu}\sin\phi))$ (B41)
$u_{0,E}^{\prime}$ $=$ $u_{0,E}+d\sin\theta_{\mu}\sin\phi$ (B42)
@@
$u_{0,N}^{\prime}$ $=$ $u_{0,N}+d\sin\theta_{\mu}\cos\phi-d\sin\theta_{s}$ (B43)
$u_{0,N}^{\prime}$ $=$ $u_{0,N}+d(\sin\theta_{\mu}\cos\phi-(\cos\phi\sin\theta_{\mu}+\cos\theta_{\mu}\sin\phi))$ (B44)
$u_{0,N}^{\prime}$ $=$ $u_{0,N}-d\cos\theta_{\mu}\sin\phi.$ (B45)
We can use Eq. B21:
$\sin\theta_{\mu}$ $=$ $\sin(\theta_{u}+90^{\circ}C)=C\cos\theta_{u}=C\hat{u}_{0,E}$ (B46)
$\cos\theta_{\mu}$ $=$ $\cos(\theta_{u}+90^{\circ}C)=-C\sin\theta_{u}=-C\hat{u}_{0,N}$ (B47)
@@
$\boxed{\boldsymbol{u_{0}}^{\prime}=\boldsymbol{u_{0}}+Cd\sin\phi\hat{u}_{0}}.$ (B50)
We may also want to go the opposite direction. To do so we can define an equivalent of Eq. B20 for $\vec{u}_{0}^{\prime}$:
$C^{\prime}\equiv(\hat{\mu}_{rel}\times\hat{u}_{0}^{\prime})\cdot\hat{z}$ (B51)
So Eqs. B46 and B47 become
$\sin\theta_{\mu}$ $=$ $\sin(\theta_{u^{\prime}}+90^{\circ}C^{\prime})=C^{\prime}\cos\theta_{u^{\prime}}=C^{\prime}\hat{u}^{\prime}_{0,E}$ (B52)
$\cos\theta_{\mu}$ $=$ $\cos(\theta_{u^{\prime}}+90^{\circ}C^{\prime})=-C^{\prime}\sin\theta_{u^{\prime}}=-C^{\prime}\hat{u}^{\prime}_{0,N}$ (B53)
@@
$u_{0,E}^{\prime}$ $=$ $u_{0,E}+C^{\prime}d\sin\phi\hat{u}^{\prime}_{0,E}$ (B54)
$u_{0,N}^{\prime}$ $=$ $u_{0,N}+C^{\prime}d\sin\phi\hat{u}^{\prime}_{0,N}$ (B55)
$\boldsymbol{u}_{0}^{\prime}$ $=$ $\boldsymbol{u}_{0}+C^{\prime}d\sin\phi\hat{u}^{\prime}_{0}$ (B56)
So Eq. B50 becomes
$\boxed{\boldsymbol{u}_{0}=\boldsymbol{u}_{0}^{\prime}-C^{\prime}d\sin\phi\hat{u}^{\prime}_{0}}.$ (B57)
B.1 Standard Coordinate Transforms
B.1.1 Between Geometric Midpoint and Primary
The separation in mas between the two lenses is $\boldsymbol{a}$ pointing towards the primary. In units of $\theta_{E}$, it’s $\boldsymbol{s}\equiv\frac{\boldsymbol{a}}{\theta_{E}}$. So the vector from the geometric midpoint to the primary is $\frac{\boldsymbol{s}}{2}$. Hence Eq. B50 becomes
$\boldsymbol{u_{\textrm{prim},0}}=\boldsymbol{u_{\textrm{geom},0}}+C\frac{a}{2\theta_{E}}\sin\phi\hat{u}_{\textrm{geom},0}.$ (B58)
When transforming from primary to geometric midpoint, Eq. B57 becomes:
$\boldsymbol{u_{\textrm{geom},0}}=\boldsymbol{u_{\textrm{prim},0}}-C^{\prime}\frac{a}{2\theta_{E}}\sin\phi\hat{u}_{\textrm{prim},0}.$ (B59)
B.1.2 Between Geometric Midpoint and Center of Mass
@@
$\boldsymbol{d}=\boldsymbol{s}\frac{1-q}{2(1+q)}\equiv\boldsymbol{s}q^{\prime}.$ (B60)
Hence Eq. B50 becomes
$\boldsymbol{u_{\textrm{com},0}}=\boldsymbol{u_{\textrm{geom},0}}+Csq^{\prime}\sin\phi\hat{u}_{\textrm{geom},0}.$ (B61)
When transforming from primary to geometric midpoint, Eq. B57 becomes:
$\boldsymbol{u_{\textrm{geom},0}}=\boldsymbol{u_{\textrm{com},0}}-C^{\prime}sq^{\prime}\sin\phi\hat{u}_{\textrm{com},0}.$ (B62)
@@
Appendix C $t_{0}$ transformation
Along with a change in the distance of closest approach, there is a change of when the closest approach occurs. In Fig. 19, the source is at $S$ at time $t_{0}$ and at $S^{\prime}$ at time $t_{0}^{\prime}$. Since the source is moving with relative proper motion $\boldsymbol{\mu_{rel}}$, we know:
$S^{\prime}-S=\frac{\boldsymbol{\mu_{rel}}}{\theta_{E}}(t_{0}^{\prime}-t_{0})$ (C1)

### v3

@@ Appendix C $u_{0}$ transformation
$\boldsymbol{u}_{0}^{\prime}$ $=$ $S^{\prime R^{\prime}}-L^{\prime R^{\prime}}=(S^{\prime R}-d\hat{s})-(L^{\prime R}-d\hat{s})=S^{\prime R}-L^{\prime R}$ (C13)
$\boldsymbol{u}_{0}^{\prime}$ $=$ $\boldsymbol{u}_{0}+\frac{\boldsymbol{\mu}_{rel}}{\theta_{E}}(t_{0}^{\prime}-t_{0})-d\hat{s}$ (C14)
@@
$\boldsymbol{\mu}_{rel}$ $=$ $[\mu_{rel}\cos\theta_{\mu},\mu_{rel}\sin\theta_{\mu}]$ (C15)
$\hat{s}$ $=$ $[\cos\theta_{s},\sin\theta_{s}]$ (C16)
Where $\theta_{\mu}$ is the angle from North to $\boldsymbol{\mu}_{rel}$ and $\theta_{s}$ is the angle from North to the binary axis, East of North. (Note that choosing North as our reference will not affect the final answer. Another reference could be chosen). The binary axis is a vector from $M_{2}$ to $M_{1}$ as indicated in Figure 18. So in components:
$u_{0,E}^{\prime}$ $=$ $u_{0,E}+\mu_{rel}\cos\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\cos\theta_{s}$ (C17)
$u_{0,N}^{\prime}$ $=$ $u_{0,N}+\mu_{rel}\sin\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\sin\theta_{s}$ (C18)
@@
$C\equiv(\hat{\mu}_{rel}\times\hat{u}_{0})\cdot\hat{z}$ (C20)
@@
$\theta_{u}=\theta_{\mu}-90^{\circ}C$ (C21)
We can now take Eqs C17 and C18 and define the components as their magnitudes times cosine/sine of angles:
$u_{0}^{\prime}\cos\theta_{u^{\prime}}$ $=$ $u_{0}\cos\theta_{u}+\mu_{rel}\cos\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\cos\theta_{s}$ (C22)
$u_{0}^{\prime}\sin\theta_{u^{\prime}}$ $=$ $u_{0}\sin\theta_{u}+\mu_{rel}\sin\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\sin\theta_{s}$ (C23)
Where $\theta_{u^{\prime}}$ is the angle from North to $\mathbf{u_{0}^{\prime}}$. $\mathbf{u_{0}}$ will always be parallel to $\mathbf{u_{0}^{\prime}}$, but they may be opposite directions. So
$\theta_{u^{\prime}}=\theta_{u}+180^{\circ}F$ (C24)
@@
$\sin\theta_{u^{\prime}}$ $=$ $G\sin\theta_{u}$ (C25)
$\cos\theta_{u^{\prime}}$ $=$ $G\cos\theta_{u}$ (C26)
where G=1G=-1 if $F=1$ and $G=1$ if $F=0$. Plugging into Eqs C22 and C23:
$u_{0}^{\prime}G-u_{0}$ $=$ $\frac{1}{\cos\theta_{u}}\Big(\mu_{rel}\cos\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\cos\theta_{s}\Big)$ (C27)
$u_{0}^{\prime}G-u_{0}$ $=$ $\frac{1}{\sin\theta_{u}}\Big(\mu_{rel}\sin\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\sin\theta_{s}\Big)$ (C28)
@@
$\cos\theta_{u}$ $=$ $\cos(\theta_{\mu}-90^{\circ}C)=C\sin\theta_{\mu}$ (C29)
$\sin\theta_{u}$ $=$ $\sin(\theta_{\mu}-90^{\circ}C)=-C\cos\theta_{\mu}$ (C30)
@@
$u_{0}^{\prime}G-u_{0}$ $=$ $\frac{1}{C}\Big(\mu_{rel}\frac{1}{\tan\theta_{\mu}}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\frac{\cos\theta_{s}}{\sin\theta_{\mu}}\Big)$ (C31)
$u_{0}^{\prime}G-u_{0}$ $=$ $-\frac{1}{C}\Big(\mu_{rel}\tan\theta_{\mu}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)-d\frac{\sin\theta_{s}}{\cos\theta_{\mu}}\Big)$ (C32)
@@
$\sin\theta_{s}$ $=$ $\sin(\theta_{\mu}+\phi)=\cos\phi\sin\theta_{\mu}+\cos\theta_{\mu}\sin\phi$ (C33)
$\cos\theta_{s}$ $=$ $\cos(\theta_{\mu}+\phi)=\cos\phi\cos\theta_{\mu}-\sin\phi\sin\theta_{\mu}$ (C34)
$\rightarrow\frac{\sin\theta_{s}}{\cos\theta_{\mu}}$ $=$ $\cos\phi\tan\theta_{\mu}+\sin\phi$ (C35)
$\frac{\cos\theta_{s}}{\sin\theta_{\mu}}$ $=$ $\frac{\cos\phi}{\tan\theta_{\mu}}-\sin\phi$ (C36)
$\rightarrow\frac{\sin\theta_{s}}{\cos\theta_{\mu}}+\frac{\cos\theta_{s}}{\sin\theta_{\mu}}$ $=$ $\cos\phi\Big(\frac{1}{\tan\theta_{\mu}}+\tan\theta_{\mu}\Big),$ (C37)
Setting Eq C31 equal to Eq C32 and simplifying:
$\mu_{rel}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)\Big(\frac{1}{\tan\theta_{\mu}}+\tan\theta_{\mu}\Big)$ $=$ $d\Big(\frac{\sin\theta_{s}}{\cos\theta_{\mu}}+\frac{\cos\theta_{s}}{\sin\theta_{\mu}}\Big)$ (C38)
$\mu_{rel}\Big(\frac{t_{0}^{\prime}-t_{0}}{\theta_{E}}\Big)$ $=$ $d\cos\phi$ (C39)
We can now use this relation in Eqs C17 and C18. Starting with the E-component:
$u_{0,E}^{\prime}$ $=$ $u_{0,E}+d\cos\theta_{\mu}\cos\phi-d\cos\theta_{s}$ (C40)
$u_{0,E}^{\prime}$ $=$ $u_{0,E}+d(\cos\theta_{\mu}\cos\phi-(\cos\phi\cos\theta_{\mu}-\sin\theta_{\mu}\sin\phi))$ (C41)
$u_{0,E}^{\prime}$ $=$ $u_{0,E}+d\sin\theta_{\mu}\sin\phi$ (C42)
@@
$u_{0,N}^{\prime}$ $=$ $u_{0,N}+d\sin\theta_{\mu}\cos\phi-d\sin\theta_{s}$ (C43)
$u_{0,N}^{\prime}$ $=$ $u_{0,N}+d(\sin\theta_{\mu}\cos\phi-(\cos\phi\sin\theta_{\mu}+\cos\theta_{\mu}\sin\phi))$ (C44)
$u_{0,N}^{\prime}$ $=$ $u_{0,N}-d\cos\theta_{\mu}\sin\phi.$ (C45)
We can use Eq. C21:
$\sin\theta_{\mu}$ $=$ $\sin(\theta_{u}+90^{\circ}C)=C\cos\theta_{u}=C\hat{u}_{0,E}$ (C46)
$\cos\theta_{\mu}$ $=$ $\cos(\theta_{u}+90^{\circ}C)=-C\sin\theta_{u}=-C\hat{u}_{0,N}$ (C47)
@@
$\boxed{\boldsymbol{u_{0}}^{\prime}=\boldsymbol{u_{0}}+C\hat{u}_{0}d\sin\phi}.$ (C50)
We may also want to go the opposite direction. To do so we can define an equivalent of Eq. C20 for $\mathbf{u_{0}^{\prime}}$:
$C^{\prime}\equiv(\hat{\mu}_{rel}\times\hat{u}_{0}^{\prime})\cdot\hat{z}$ (C51)
So Eqs. C46 and C47 become
$\sin\theta_{\mu}$ $=$ $\sin(\theta_{u^{\prime}}+90^{\circ}C^{\prime})=C^{\prime}\cos\theta_{u^{\prime}}=C^{\prime}\hat{u}^{\prime}_{0,E}$ (C52)
$\cos\theta_{\mu}$ $=$ $\cos(\theta_{u^{\prime}}+90^{\circ}C^{\prime})=-C^{\prime}\sin\theta_{u^{\prime}}=-C^{\prime}\hat{u}^{\prime}_{0,N}$ (C53)
@@
$u_{0,E}^{\prime}$ $=$ $u_{0,E}+C^{\prime}\hat{u}^{\prime}_{0,E}d\sin\phi$ (C54)
$u_{0,N}^{\prime}$ $=$ $u_{0,N}+C^{\prime}\hat{u}^{\prime}_{0,N}d\sin\phi$ (C55)
$\boldsymbol{u}_{0}^{\prime}$ $=$ $\boldsymbol{u}_{0}+C^{\prime}\hat{u}^{\prime}_{0}d\sin\phi$ (C56)
So Eq. C50 becomes
$\boxed{\boldsymbol{u}_{0}=\boldsymbol{u}_{0}^{\prime}-C^{\prime}\hat{u}^{\prime}_{0}d\sin\phi}.$ (C57)
C.1 Standard Coordinate Transforms
C.1.1 Between Geometric Midpoint and Primary
The separation in mas between the two lenses is $\boldsymbol{a}$ pointing towards the primary. In units of $\theta_{E}$, it’s $\boldsymbol{s}\equiv\frac{\boldsymbol{a}}{\theta_{E}}$. So the vector from the geometric midpoint to the primary is $\frac{\boldsymbol{s}}{2}$. Hence Eq. C50 becomes
$\boldsymbol{u_{\textrm{prim},0}}=\boldsymbol{u_{\textrm{geom},0}}+C\hat{u}_{\textrm{geom},0}\frac{a}{2\theta_{E}}\sin\phi.$ (C58)
When transforming from primary to geometric midpoint, Eq. C57 becomes:
$\boldsymbol{u_{\textrm{geom},0}}=\boldsymbol{u_{\textrm{prim},0}}-C^{\prime}\hat{u}_{\textrm{prim},0}\frac{a}{2\theta_{E}}\sin\phi.$ (C59)
C.1.2 Between Geometric Midpoint and Center of Mass
@@
$\boldsymbol{d}=\boldsymbol{s}\frac{1-q}{2(1+q)}\equiv\boldsymbol{s}q^{\prime}.$ (C60)
Hence Eq. C50 becomes
$\boldsymbol{u_{\textrm{com},0}}=\boldsymbol{u_{\textrm{geom},0}}+C\hat{u}_{\textrm{geom},0}sq^{\prime}\sin\phi.$ (C61)
When transforming from primary to geometric midpoint, Eq. C57 becomes:
$\boldsymbol{u_{\textrm{geom},0}}=\boldsymbol{u_{\textrm{com},0}}-C^{\prime}\hat{u}_{\textrm{com},0}sq^{\prime}\sin\phi.$ (C62)
@@
Appendix D $t_{0}$ transformation
Along with a change in the distance of closest approach, there is a change of when the closest approach occurs. In Figure 19, the source is at $S$ at time $t_{0}$ and at $S^{\prime}$ at time $t_{0}^{\prime}$. Since the source is moving with relative proper motion $\boldsymbol{\mu_{rel}}$, we know:
$S^{\prime}-S=\frac{\boldsymbol{\mu_{rel}}}{\theta_{E}}(t_{0}^{\prime}-t_{0})$ (D1)